Your firm needs a computerized machine tool lathe which costs $41,000 and requires $11,100 in maintenance for each year of its 3-year life. After three years this machine will be replaced. The machine falls into the MACRS 3-year class life category and neither bonus depreciation nor Section 179 expensing can be used. Assume a tac rate of 21 percent and a discount rate of 11 percent.

If the lathe can be sold for $4,100 at the end of year 3 what is the after tax salvage value?
Answer #1

rate positively ..

Computation of after tax salvage value
MARCS depreciation table as below
Year 3-Year
1 33.33%
2 44.45%
3 14.81%
4 7.41%
Book value at year 3 =
Cost * depreciation rate for 4th year
41000*7.41% 3038.10
i Sales price = 4100
ii Less Book value = 3038.10
iii=i-ii Gain on sale 1061.90
iv=iii*21 Tax on gain @ 21% 222.999
v=i-iv Post tax salvage value= 3877.0
ans = 3877.0